# Feedback Friction: LLMs Struggle to Accurately Incorporate External Feedback

### v1

@@ 6 Limitations
While our study successfully demonstrates the existence of Feedback Friction and rules out several potential causes (model confidence, data familiarity, reasoning complexity), we lack a definitive mechanistic explanation for why models resist incorporating feedback. A better understanding of Feedback Friction likely involves complex interactions between feedback understanding, instruction following, and belief updating. Future work would benefit from more sophisticated mechanistic interpretability techniques, such as causal intervention methods and circuit analysis [64, 65], to understand the specific computational pathways through which feedback resistance emerges and persists across model architectures.
@@
Despite extensive experimentation with sampling strategies, our attempts to fully mitigate feedback friction yielded only modest improvements. The most promising next step appears to be supervised fine-tuning or reinforcement learning approaches that could (1) enhance the solver model’s receptiveness to feedback incorporation, and (2) improve the feedback generator’s ability to provide more effective guidance [66]. However, the computational constraints of our experimental setup, particularly the large model sizes we tested (including 70B+ parameter models like Claude 3.7 and Llama-4-Maverick) and the associated computational costs, prevented us from conducting comprehensive fine-tuning experiments that might meaningfully address feedback resistance.
@@ Acknowledgment
This work is supported by ONR grant (N00014-241-2089). The GPUs were provided by the DSAI and ARCH clusters. We sincerely thank the broader JHU community for discussions and feedback.
@@ Appendix D Analysis of model confidence and Feedback Friction
Initial $\text{Initial Confidence}=\exp\left(\frac{1}{T}\sum_{t=1}^{T}\log p(a_{t}\mid a_{<t},q)\right)$

### v2

@@ 6 Limitations
While our study identifies key insights into Feedback Friction—including the correlation between semantic entropy and feedback resistance and the disconnect between models’ stated understanding and actual behavior, we lack a definitive mechanistic explanation for why models resist incorporating feedback. A better understanding of Feedback Friction likely involves complex interactions between feedback understanding, instruction following, and belief updating. Future work would benefit from more sophisticated mechanistic interpretability techniques, such as causal intervention methods and circuit analysis [65, 66], to understand the specific computational pathways through which feedback resistance emerges and persists across model architectures.
@@
Despite extensive experimentation with sampling strategies, our attempts to fully mitigate feedback friction yields only modest improvements. The most promising next step appears to be supervised fine-tuning or reinforcement learning approaches that could (1) enhance the solver model’s receptiveness to feedback incorporation, and (2) improve the feedback generator’s ability to provide more effective guidance [67]. However, the computational constraints of our experimental setup, particularly the large model sizes we tested (including 70B+ parameter models like Claude 3.7 and Llama-4-Maverick) and the associated computational costs, prevented us from conducting fine-tuning experiments that might meaningfully address feedback resistance.
@@ Acknowledgment
This work is supported by ONR grant (N00014-241-2089). The GPUs were provided by the DSAI and ARCH clusters. We sincerely thank Yuqing Yang, Zhengping Jiang, and the broader JHU community for discussions and feedback.
@@ Appendix D Analysis of model confidence and Feedback Friction
$\text{Initial Confidence}=\exp\left(\frac{1}{T}\sum_{t=1}^{T}\log p(a_{t}\mid a_{<t},q)\right)$
@@ Appendix G Analysis of model type and Feedback Friction
Appendix H Details of testing model’s self-perception of Feedback Friction
Figure 19 shows the prompts we used to probe whether the model understand the given feedback and whether it’s willing to change its belief.
Figure 19: An example of Llama-4-Scout explicitly acknowledging changing its belief while still making the same mistake in the next attempt.
Figure 20 shows an example where Llama-4-Scout explicitly acknowledges that “the actual boundary is defined by the perpendicular bisectors of the segments connecting O to the vertices, which can result in an oblique line” and claims its approach has been updated. Yet in its second attempt, it still describes the region as forming “a square with side length 1/2 min(AB, AD)” and arrives at the same incorrect answer of 1/4. This demonstrates the model can articulate understanding of the feedback while its actual problem-solving behavior remains unchanged. The correct answer is 1/2.
Figure 20: An example of Llama-4-Scout explicitly acknowledging changing its belief while still making the same mistake in the next attempt.
Appendix I Full results of semantic entropy
Figure 21 presents the complete self-correction analysis across all five benchmark tasks (5 digits multiplication, GPQA, MATH, MMLU-Pro, and TriviaQA) for both Llama-3.3-70B and Llama-4-Scout. Each subplot shows initial accuracy (blue), final accuracy (orange), and absolute improvement rate (green) as functions of semantic entropy with bucket size 0.2. Regions with fewer than 10 samples are marked with red shading and dashed boundaries to indicate unreliable data, with sample counts displayed in boxes above each plot.
Figure 21: Full results of semantic entropy over five benchmark tasks and two models.